### v1

# Cristiá, Capozucca and Rossi

Abstract
$\{log\}$ (read ‘setlog’) was born as a Constraint Logic Programming (CLP) language where sets and binary relations are first-class citizens, thus fostering set programming. Internally, $\{log\}$ is a constraint satisfiability solver implementing decision procedures for several fragments of set theory. Hence, $\{log\}$ can be used as a declarative, set, logic programming language and as an automated theorem prover for set theory. Over time $\{log\}$ has been extended with some components integrated to the satisfiability solver thus providing a formal verification environment. In this paper we make a comprehensive presentation of this environment which includes a language for the description of state machines based on set theory, an interactive environment for the execution of functional scenarios over state machines, a generator of verification conditions for state machines, automated verification of state machines, and test case generation. State machines are both, programs and specifications; exactly the same code works as a program and as its specification. In this way, with a few additions, a CLP language turned into a seamlessly integrated programming and automated proof system.
keywords:
$\{log\}$, set theory, constraint logic programming, formal verification
1 Introduction
The work presented in this paper is aligned with the following quotations (Garavel et al.,, 2020).
The introduction of programming languages supported by proof systems. I do expect it to become more common with programming language implementations being born with program verifiers of various kinds. […] It is a pleasure to see modern programming languages to an increasing degree look and feel like well known formal specification languages.
Another challenge is to better integrate specification and proof with programming, preferably at the level of programming languages and tools.
That is, this paper presents $\{log\}$ (read ‘setlog’) (Rossi and Cristiá,, 2000; Dovier et al.,, 2000), a programming language that has the look-and-feel of a formal specification language which in turn is supported by a proof system. The core of $\{log\}$ is a Constraint Logic Programming (CLP) language and interpreter written in SWI-Prolog implementing a constraint satisfiability solver for set theory. The satisfiability solver is a rewriting system implementing decision procedures for several fragments of set theory (e.g., integer intervals (Cristiá and Rossi, 2024c, ) and restricted quantifiers (Cristiá and Rossi, 2024d, )). In this way, $\{log\}$ implements declarative programming in the form of set programming Schwartz et al., (1986); Cantone et al., (2001). As such $\{log\}$ has been applied to several case studies showing that the satisfiability solver works in practice (Cristiá and Rossi, 2021a, ; Cristiá and Rossi, 2021c, ; Cristiá and Rossi, 2024a, ; Cristiá et al.,, 2023; Capozucca et al.,, 2024).
However, writing good $\{log\}$ code, adding the corresponding verification conditions and passing them to the satisfiability solver as to ensure the code verifies some properties, can be cumbersome. In this context, good $\{log\}$ code means code for which it is easy to write down verification conditions that can be automatically discharged by $\{log\}$. As a programming language, $\{log\}$ includes many programming facilities leading to code that is not always easy to verify because, many times, lies outside the fragments of set theory for which $\{log\}$ implements decision procedures. Hence, if a verification condition is given by a $\{log\}$ formula for which no decision procedure can be applied, $\{log\}$ will not be able to automatically discharge it. This, in turn, implies less reliable code that needs more human effort to be verified.
For these reasons, over the years, $\{log\}$ has been extended with some components (also implemented in SWI-Prolog) that constrain programmers to write code that is easier to verify. This programming style within $\{log\}$ defines a language for the description of state machines in terms of first-order logic (FOL), set theory and integer arithmetic—much as in formal notations such as B (Abrial,, 1996) and Z (Woodcock and Davies,, 1996). In this way: (i) the state machine language allows for the automatic generation of verification conditions (VC); (ii) these VC are automatically fed into the satisfiability solver which informs the user if they could be discharged or not (in which case, a counterexample is generated); (iii) state machines can still be executed as regular $\{log\}$ code thus allowing users to use them as working functional prototypes before attempting any serious proof campaign; and (iv) users can generate test cases to test a future, more efficient implementation written in some other programming language. This paper makes a comprehensive presentation of these extensions aiming at showing how theoretical contributions in the fields of CLP and set programming can be turned into a feasible tool aimed at formal verification.
As the above quotations suggest, there are other systems where programming and proof coexist. Four conspicuous, widely known, successful representatives of this kind are Agda (Logic and Types group,, 2004), Dafny (Leino,, 2010), F* (read ‘f-star’) (Microsoft Research and INRIA,, 2015) and Why3 (Toccata group,, 2012). All of them have been applied to a range of problems, some times of industrial size and complexity. Agda is a typed functional programming language and proof assistant based on intuitionistic type theory. The functional language provides usual functional features such as inductive types, pattern matching, and so on. Proofs are interactive and written in a functional programming style that, unlike Coq (now renamed as Rocq) (Bertot and Castéran,, 2004), does not provide proof tactics. Programmers can, nonetheless, implement Agda functions that return a proof of some property [^1] Writing such proofs requires an unusual degree of training in formal software development.. These functions embody sort of proof tactics. Later, these functions are run during type checking: if one of them fails, then type checking fails. This means that the program does not verify a certain property. As a form of proof automation Agda provides proof search by enumerating possible proof terms that may fit in the specification. Dafny combines ideas from the functional and imperative paradigms, including some support for object-oriented programming. Programs are annotated with specifications through preconditions, postconditions, loop invariants, etc. The specification language has the look-and-feel of a programming language including mathematical integers, bit-vectors, sequences, sets, induction, and so on. Proof obligations are automatically discharged, given sufficient specification, by the Z3 SMT solver (Microsoft Research,, 2008). Sufficient specification means, in general, to write sufficient proof steps (e.g., the inductive step in an inductive proof) the remainder of which are discharged automatically. F* supports purely functional programming as well as effectful programming (McBride and Paterson,, 2008). Its powerful (dependent) type system allows to describe precise program specifications including security and functional properties. The typechecker will check whether or not the program verifies its specification by discharging proofs interactively written by the programmer (using tactics, metaprogramming and symbolic computation) or by calling the Z3 SMT solver. F* shares some features with Agda. Why3 is a platform for deductive program verification. It provides WhyML, a specification and programming language. Why3 uses many external theorem provers (e.g., Alt-Ergo (OCamlPro,, 2006), CVC4-5 (Barrett et al.,, 2011; Barbosa et al.,, 2022) and E-prover (Schulz,, 1999; Schulz et al.,, 2019)) to discharge proof obligations. The system includes a library of mathematical theories (e.g., arithmetic and sets) and some programming data structures (e.g., arrays and hash tables).
Having a different origin than the systems mentioned above and not being strictly programming and proof systems, we should also mention Atelier-B (Clearsy,, 2009) and Rodin (Dep. Sys. and Soft. Eng. Grp.,, 2004). These tools are meant to be used to write and verify B and Event-B (Abrial,, 2010) specifications, respectively. In both languages, specifications take the form of state machines described by means of FOL and set theory. This is the main point of contact with $\{log\}$. Both tools rely on automatic and interactive provers, including internal ones and external SMT solvers such as VeriT (Déharbe and Fontaine,, 2009) and CVC4. B and Event-B specifications can be executed (or animated) if the ProB tool (Leuschel,, 2003) is available in Atelier-B and Rodin. ProB is a constraint solver developed in Prolog. Unlike $\{log\}$, in general, ProB cannot prove properties although it can disprove them.
Contributions and novelty.
The main contribution of this paper is a comprehensive presentation of the following $\{log\}$ extensions:
  1. 1.
    A declarative state machine specification language defined on top of the $\{log\}$ constraint language. State machines are specified in terms of set theory and FOL.
  2. 2.
    The Next environment and other facilities to simplify the execution and analysis of functional scenarios based on state machines.
  3. 3.
    A verification condition generator (VCG) that generates a set of standard VC for state machines. VC are discharged by calling $\{log\}$ itself, and if a VC could not be discharged, it allows users to analyze why this happened.
  4. 4.
    An implementation of a model-based testing framework for $\{log\}$ state machines that generates test cases by calling $\{log\}$ itself. Users can generate test cases from the state machine to test its implementation.
None of these extensions have been published before nor they have been presented as an integrated verification framework with $\{log\}$ at its center.
Concerning novelty, $\{log\}$ is quite different from all the systems mentioned above: (a) it provides a single language for properties, specifications and programs, there is no distinction between programs and specifications, the same code works as a program and its specification; (b) the same solver is used to execute programs, to prove their properties and to generate test cases, no external solvers are needed; (c) sets and binary relations are first-class entities thus naturally raising the abstraction level of specifications; and (d) all proofs are automatic as the result of implementing decision procedures for several expressive fragments of set theory.
Although each of these characteristics are not new in the CLP community, as far as we know, no other tool presents all of them together in the context of formal verification. We believe $\{log\}$ provides evidence that CLP techniques are as valuable for formal verification as techniques coming from the functional programming realm.
Structure of the paper.
The paper starts by introducing $\{log\}$ in a user-friendly manner in Section 2. Section 3 describes the state machine specification language by means of a running example. These state machines can be used as functional prototypes. We show how to run and analyze functional scenarios over these prototypes in Section 4. The VCG is described in Section 5. This section includes how $\{log\}$ can be used as an automated theorem prover, how to analyze undischarged verification conditions, and an informal account of the classes of formulas fitting in the decision procedures implemented in $\{log\}$. Section 6 describes the model-based testing method implemented in $\{log\}$. We close the paper with some concluding remarks in Section 7.
2 The $\{log\}$ CLP language and satisfiability solver
$\{log\}$ is a publicly available constraint satisfiability solver and a declarative set-based, constraint-based programming language implemented in SWI-Prolog (Rossi and Cristiá,, 2000; Dovier et al.,, 2000). $\{log\}$ is deeply rooted in the work on Computable Set Theory (Cantone et al.,, 1989), combined with the ideas put forward by the set-based programming language SETL (Schwartz et al.,, 1986).
$\{log\}$ implements various decision procedures for different theories on the domain of finite sets and integer numbers. Specifically, $\{log\}$ implements decision procedures for: (1) The theory of hereditarily finite sets, i.e., finitely nested sets that are finite at each level of nesting (Dovier et al.,, 2000); (2) A very expressive fragment of the theory of finite set relation algebras (Cristiá and Rossi,, 2020, 2018); (3) Theory (1) extended with restricted intensional sets (Cristiá and Rossi, 2021b, ); (4) Theory (1) extended with cardinality constraints (Cristiá and Rossi,, 2023); (5) Theory (4) extended with integer intervals (Cristiá and Rossi, 2024c, ); (6) Quantifier-free, decidable languages extended with restricted quantifiers (Cristiá and Rossi, 2024d, ); and (7) The theory of linear integer arithmetic, by integrating an existing decision procedure for this theory. All these procedures constitute the core of the $\{log\}$ tool. Several in-depth empirical evaluations provide evidence that $\{log\}$ is able to solve non-trivial problems (Cristiá and Rossi,, 2020, 2018; Cristiá and Rossi, 2021b, ; Cristiá et al.,, 2013); in particular as an automated verifier of security properties (Cristiá and Rossi, 2021a, ; Cristiá and Rossi, 2021c, ; Cristiá et al.,, 2023; Capozucca et al.,, 2024).
Remark 1 (Scope of presentation)
Below we introduce $\{log\}$. The presentation is brief and user-oriented. A complete, user-oriented presentation can be found in the $\{log\}$ user’s manual (Rossi and Cristiá,, 2025). A formal presentation can be found in the papers cited in the previous paragraph, including formal syntax and semantics; and soundness, completeness and termination proofs. We assume a basic knowledge of (constraint) logic programming and Prolog.
In the introduction we say that in $\{log\}$ there is no distinction between programs and specifications. Let’s say we need a program computing the minimum element of a set of integer numbers. The specification of that program can be the following predicate where $S$ is a set and $m$ its minimum element ($\triangleq$ means equal by definition):
$smin(S,m)\triangleq m\in S\land(\forall x\in S:m\leq x)$ (1)
The $\{log\}$ implementation of $smin$ is the following predicate:
smin(S,M) :- M in S & foreach(X in S, M =< X). (2)Right at first glance the differences between specification and program look minimal. For instance, we write M instead of $m$ to denote a variable (as in Prolog), in denotes set membership ($\in$), & denotes conjunction ($\land$), and foreach denotes the restricted universal quantifier ($\forall x\in S:\dots$). Indeed, the following is another implementation for $smin$:
smin1(S,M) :- foreach(X in S, M =< X) & M in S.That is, & enjoys commutativity as much as $\land$. In other words, both smin and smin1 will produce the same results—although they can have different performance. The commutativity of & makes smin look as a formula. But it is also a program as we can call it to get the minimum element of a set:
{log}=> smin({9,-3,5,1,8},M).  M = -3Hence, smin is called with an input set and $\{log\}$ computes its minimum (M = -3). Furthermore, the input set can be partially specified (i.e., some elements can be variables) [^2] The computed answers are simplified for presentation purposes.:
{log}=> smin({9,-3,5,X,8},M).    % variable X in the set  M = -3  Constraint: -3 =< X  ;  M = X  Constraint: X =< 9, X =< -3, X =< 5, X =< 8In $\{log\}$ a solution is a (possibly empty) list of equalities of the form $var=term$ and a (possibly empty) list of constraints. The list of constraints is guaranteed to be always satisfiable. In fact we can see a possible value for X if we ask $\{log\}$ to produce ground solutions (i.e., solutions where variables occur only at the left-hand side of bindings):
{log}=> groundsol.{log}=> smin({9,-3,5,X,8},M).  X = -3,  M = -3
As in Prolog, $\{log\}$ does not distinguish between inputs and outputs. The fact that so far we have considered that S is an input and M is an output is purely out of convenience and common use. $\{log\}$ can compute the other way round:
1{log}=> smin(S,10).2  S = {10 / _N1}3  Constraint: set(_N1), foreach(_X1 in _N1, 10 =< _X1)4{log}=> groundsol.5{log}=> smin(S,10).6  S = {10}In the first run (line 1), $\{log\}$ states that for 10 to be the minimum of S, then S must be a set of the form {10/_N1} meaning that 10 belongs to S and that all of its other possible elements (represented by set _N1) are grater than or equal to 10. Again, the constraint is always satisfiable—in this case by substituting _N1 by the empty set, which is the solution computed after the activation of groundsol (lines 4-6). The term {10/_N1} corresponds to $\{10\}\cup\_N1$ what implies that {10/{}} equals {10}.
smin is not a formula only because of the commutativity of & but, fundamentally, because it is possible to prove properties true of it. The following is a known property of the minimum of a set:
$smin(A,m)\land smin(B,n)\land A\subseteq B\implies n\leq m$ (3)
Given that $\{log\}$ is a satisfiability solver, in order to check whether or not smin verifies this property we have to ask $\{log\}$ if the negation of the property is unsatisfiable. So we run:
{log}=> neg(smin(A,M) & smin(B,N) & subset(A,B) implies N =< M).  noAs $\{log\}$ answers no we know the formula is unsatisfiable—i.e., there are no finite sets A and B and integers M and N that can satisfy the formula. Therefore, the formula inside neg is always satisfiable or valid.
Hence, we can use smin as both a formula or specification and as a program or implementation. The same piece of code, no need for text describing the specification and another text describing the implementation, no need for program annotations. We call this the program-formula duality or the implementation-specification duality. Some times we refer to $\{log\}$ code as forgrams, a portmanteau word resulting from the combination of formula and program. Therefore, when $\{log\}$ programmers write code they are writing a program as a formula (Apt and Bezem,, 1999). The same solver executes a program and proves properties true of it, no need for external solvers or theorem provers. Actually, $\{log\}$ cannot be divided into a component running programs and a component proving properties.
Although some of these ideas are not new to the CLP community, it is our understanding that no other tool features set programming based on decision procedures implemented with CLP properties—e.g. the program-formula duality.
2.1 Extensional sets, relations and functions
[…]
{log}=> un({X/A},{3,a/B},{Y/C}).  Y = X,  C = {3,a/_N1}  Constraint: X nin A, un(A,B,_N1), X neq 3, X neq a, X nin _N1where X nin A denotes $X\notin A$ and X neq 3 denotes $X\neq 3$—note that a is a constant, not a variable.
The constraint size(A,N) denotes $|A|=N$, i.e., set cardinality. The first argument can be an extensional set or the empty set; the second one can only be a variable or an integer number, but it can participate in integer constraints. Hence, $\{log\}$ can compute the solutions (or lack thereof) of444is is the standard Prolog predicate for the evaluation of integer expressions.:
{log}=> groundsol.{log}=> size(A,N) & size(B,M) & B neq {} & 5 is M + N & 2 =< N - M.  A = {n0,n1,n2,n3},  N = 4,  B = {n4},  M = 1
In $\{log\}$ a binary relation is just a set of ordered pairs. The ordered pair $(x,y)$ is written [x,y]. $\{log\}$ provides all the operators of set relation algebra in the form of constraints. Then, for example, comp(R,S,T) denotes composition of binary relations ($T=R;S$) and inv(R,S) denotes the converse (or inverse) of a binary relation $(S=R^{\smile}$). Cartesian product is also available: cp(A,B) denotes the set $A\times B$. Extensional sets and the empty set can be arguments to all relational operators and Cartesian product.
[…]
Functions are a subclass of binary relations—i.e., functions are sets of ordered pairs. The constraint pfun(F) states that the binary relation F is a function. $\{log\}$ provides two notions of function application: if pfun(F) holds then apply(F,X,Y) is satisfied iff $X\in dom(F)\land F(X)=Y$; in turn, applyTo(F,X,Y) denotes $\{(X,X)\};F=\{(X,Y)\}$, for any binary relation $F$.
Given that relations and functions are sets, they can be combined and passed as arguments to all the classic set and relational operators.
2.2 More set terms
Two more set terms are available in $\{log\}$. Terms of the form int(m,n), with both arguments either variables or integer numbers, denote integer intervals—i.e., sets of the form $[m,n]\cap\mathbb{Z}$. Interval limits can participate in integer constraints. Intervals can be arguments to all the classic set operators including cardinality—but in general they cannot be passed as arguments to relational operators. In this way $\{log\}$ is able to compute solutions (or lack thereof) for goals such as:
{log}=> groundsol.{log}=> un({X,Y},{Z/A},int(5,M)) & X neq Y & 7 < M.  X = 5,  Y = 6,  Z = 7,  A = {8},  M = 8
Finally, $\{log\}$ provides a set term, called Restricted Intensional Set (RIS), denoting set comprehensions of the form $\{x\in A:\phi(x)\}$ where $\phi$ is some formula depending on $x$. For instance, the set $\{x\in A\mid 3x+2\geq 0\}$ is encoded in $\{log\}$ as the term ris(X in A, 3*X + 2 >= 0), where A can be an extensional set term or another RIS. RIS have a more expressive and complex structure that we are not going to show here. RIS terms can only be used as arguments of classic set operators not including cardinality—they cannot be passed as arguments to relational operators.
2.3 Negation
Negation in $\{log\}$ is a delicate matter as it is in logic programming in general (Apt and Bol,, 1994). The problem with negating a $\{log\}$ formula is that its result may not be a $\{log\}$ formula. Consider the following alternative implementation of $smin$ (1):
smin2(S,M) :- M in S & subset(S,int(M,Max)).Note the presence of Max, an unbound variable making part of the body but not of the predicate’s head. This is an existential variable. Thus, the negation of smin2 requires the introduction of a (unrestricted) universal quantification over Max, that is:
$\forall Max(\lnot(M\in S\land S\subseteq[M,Max])).$ (4)
The problem is that $\{log\}$ does not admit (unrestricted) universal quantification, meaning that (4), as it is, cannot be encoded as a $\{log\}$ formula. That does not mean that there is no $\{log\}$ formula encoding the negation of smin2. Actually, the following is such a formula:
$M\notin S\lor X\in S\land X<M$ (5)
were $X$ is an existential variable. The point is that, in general, neg(smin2) cannot be automatically computed. Fortunately, in the case of $smin$, since its encoding with smin in (1) contains no existential variables, $\{log\}$ is able to automatically compute neg(smin):
$\lnot(M\in S\land\forall X\in S:M\leq X)\equiv M\notin S\lor\exists X\in S:X<M$ (6)
where the right-hand side predicate is encoded in $\{log\}$ as:
M nin S or exists(X in S, X < M).
Now consider the following $\{log\}$ predicate:
property(R,S) :- dom(R,Dr) & dom(S,Ds) & subset(Dr,Ds).where dom(R,D) encodes the domain of relation R, i.e., $dom(R)=D$. Clearly, property introduces two existential variables (Dr and Ds) thus making it impossible to use neg to compute its negation. However, the nature of these variables is different from Max in the $smin$ example. Here, these variables define the domain of R and S. In other words, Dr and Ds are sort of names for the expressions $dom(R)$ and $dom(S)$. This means that the negation of property is not:
$\forall Dr,Ds(dom(R)\neq Dr\lor dom(S)\neq Ds\lor Dr\not\subseteq Ds)$ (7)
It makes no sense to state $dom(R)\neq Dr$ for all $Dr$ because it would mean that the domain of $R$ does not exist. For cases like this $\{log\}$ provides the Let predicate (Cristiá and Rossi, 2024d, , Sect. 5):
property(R,S) :- let([Dr,Ds], dom(R,Dr) & dom(S,Ds), subset(Dr,Ds)).Let is interpreted as follows:
[…]
$\triangleq\exists x_{1},\dots,x_{n}(x_{1}=e_{1}\land\dots\land x_{n}=e_{n}\land\phi)$
$\equiv\phi[e_{1}/x_{1},\dots,e_{n}/x_{n}]$ [by one-point rule (Woodcock and Davies,, 1996, Sect 4.2)]
where $x_{i}$ are variables, $e_{i}$ terms such that $x_{i}$ does not occur in $e_{i}$ and $\phi$ is a formula. Therefore, negating property(R,S) proceeds as follows:
$\lnot property(R,S)$
[…]
$\equiv\lnot dom(R)\subseteq dom(S)$
where the rightmost formula can be encoded in $\{log\}$ as follows:
dom(R,Dr) & dom(S,Ds) & neg(subset(Dr,Ds))
The Let predicate allows us to use neg for a whole class of formulas for which, otherwise, negation would have needed human intervention. In summary, neg($P$) is safe provided $P$ does not contain existential variables.
2.4 Restricted quantifiers
We have already introduced restricted quantifiers (RQ) in $\{log\}$. We have seen foreach (restricted universal quantifier, RUQ) and exists (restricted existential quantifier, REQ), in the previous sections. Here we discuss them a little more deeply. First of all, RQ in $\{log\}$ are at least as expressive as set relation algebra (Cristiá and Rossi, 2024d, , Sect. 4.3), in part, due to the following characteristics. First, RQ admit quantification terms rather than simply variables. For example:
foreach([X,Y] in R, X is Y + 1)states that all the elements of R of the form [X,Y] must verify X is Y + 1. As [X,Y] denotes an ordered pair, then R is expected to be a binary relation.
Second, RQ can be nested. For instance:
foreach([[X,Y] in R, [V,W] in S], X = V implies Y is W + 2)is equivalent to the following nested RUQ:
foreach([X,Y] in R, foreach([V,W] in S, X = V implies Y is W + 2))Third, a bound variable can be used as quantification domain in an inner RQ:
foreach([X,Y] in R, exists(V in Y, X is V + 1))  (8)That is, the elements of the range of R are expected to be sets—due to the set membership constraint V in Y.
The introduction of REQ in the language deserves a special note. Since logic programming allows for the introduction of existential variables, the presence of REQ might seem unnecessary, but it is not. REQ help in further extending the fragment of the language where negation can be easily computed. Consider the following simple predicate:
p(S) :- X in S & X > 0.The negation of p cannot be easily computed due to the presence of X, an existential variable. However, p can be written in terms of a REQ preserving its meaning while exchanging an existential variable for a bound one:
p(S) :- exists(X in S, X > 0).Now neg(p(S)) can be easily computed:
$\lnot p(S)\equiv\lnot(\exists X\in S:X>0)\equiv(\forall X\in S:\lnot(X>0))\equiv(\forall X\in S:X\leq 0)$ (9)
where the rightmost predicate is written in $\{log\}$ as: foreach(X in S, X =< 0).
2.5 Types
Rooted in Prolog, $\{log\}$ provides essentially an untyped language based on untyped FOL and untyped set theory. In this way, a set such as $\{1,a,(3,b),\{x,q\}\}$ is perfectly legal in $\{log\}$. Untyped formalisms are not bad in themselves but types help to avoid some classes of errors (Lamport and Paulson,, 1999). For this reason, we recently defined a type system and implemented a typechecker as a $\{log\}$ component (Cristiá and Rossi, 2024b, ).
Users can activate/deactivate the typechecker according to their needs. The type system defines basic or uninterpreted types, types for integers and strings, sum types [^5] Tagged union, variant, variant record, choice type, discriminated union, disjoint union, or coproduct., Cartesian products, and set types—in this sense, the type system is similar to those used in B and Z. Set and relational operators are polymorphic. Users must declare the type of user-defined predicates. We will see more about types in Section 3.
3 A logic language for state machines
As we have explained in the introduction, we have defined a language on top of $\{log\}$ for the description of state machines. The language is inspired in the B and Z specification languages. This language constrains programmers to use $\{log\}$ in a more restricted way but increasing the chances of automatic verification. We will introduce the main elements of the language by means of a running example. Some elements are deliberately left out of the paper for brevity—see $\{log\}$ user’s manual for more details (Rossi and Cristiá,, 2025, Section 13.1).
The specification to be used as running example is known as the birthday book. It is a system which records people’s birthdays, and is able to issue a reminder when the day comes round. The problem is borrowed from Spivey (1992).
A $\{log\}$ state machine is a collection of declarations and predicates. Predicates and declarations are similar to Prolog’s. State machines transition between states defined by means of state variables. The first declaration is used to declare the state variables of the state machine. In the birthday book we have the following declaration:
variables([Known, Birthday]).where Known is intended to be the set of names with birthdays recorded; and Birthday is a function which, when applied to certain names, gives the birthdays associated with them.
After introducing the state variables, one or more state invariants can be declared. In the birthday book we have the following two:
invariant(domBirthday).domBirthday(Known,Birthday) :- dom(Birthday,Known).
invariant(birthdayFun).birthdayFun(Birthday) :- pfun(Birthday).As can be seen, each invariant is preceded by an invariant declaration. The body of each invariant is a $\{log\}$ formula depending on the state variables. An invariant declaration is a declaration of intent. It remains to be proved that these predicates are indeed preserved by the operations of the specification—see Section 5.
In the state machine language recursion is not allowed—although $\{log\}$ admits recursive predicates. Hence, invariants, as well as all the other elements of the language, have to be given as non-recursive $\{log\}$ predicates. In spite that this can be seen as a severe restriction, it is not. The main elements of specification languages such as Z and B do not admit recursive definitions. Given the applicability of Z and B, the lack of recursion in the $\{log\}$ state machine language should not be a hard limitation.
Once all the invariants have been given, the set of initial states of the specification has to be given. In many specifications a single initial state is given. This is the case of the birthday book.
initial(bbInit).bbInit(Known,Birthday) :- Known = {} & Birthday = {}.
[…]
addBirthday(Known,Birthday,Name,Date,Known_,Birthday_)where Name and Date are the inputs; and Known and Birthday represent the before-state while Known_ and Birthday_ represent the after-state. We will define addBirthday by splitting it in a couple of predicates. The first one specifies the case when the given name and date can actually be added to the system, i.e., when the name is new to the system:
addBirthdayOk(Known,Birthday,Name,Date,Known_,Birthday_) :-  Name nin Known &  un(Known,{Name},Known_) &  un(Birthday,{[Name,Date]},Birthday_).It is easy to see that the specification is given by providing pre and postconditions. For instance, the constraint Name nin Known is a precondition whereas un(Birthday,{[Name,Date]},Birthday_) is a postcondition stating that the value of Birthday in the next state is equal to the union of its value in the before state with {[Name,Date]}.
The second predicate describes what to do if the user wants to add a name already present in the system:
nameAlreadyExists(Known,Birthday,Name,Known_,Birthday_) :-  Name in Known & Known_ = Known &  Birthday_ = Birthday.That is, the system has to remain in the same state. Finally, we declare the full operation:
operation(addBirthday).addBirthday(Known,Birthday,Name,Date,Known_,Birthday_) :-  addBirthdayOk(Known,Birthday,Name,Date,Known_,Birthday_)  or  nameAlreadyExists(Known,Birthday,Name,Known_,Birthday_).In this case we precede the clause by the operation declaration. Note that addBirthdayOk and nameAlreadyExists are not operations, although they participate in one.
Now we specify an operation listing all the persons whose birthday is a given date:
operation(remind).remind(Birthday,Today,Names) :-  rres(Birthday,{Today},M) & dom(M,Names).where rres(R,B,S) is a constraint called range restriction whose interpretation is $S=\{(x,y)\in R\mid y\in B\}$777rres/3 is one of the relational operators that can be defined in terms of the classic set operators and the basic operators of set relation algebra.. Note that state variable Known is not an argument simply because the operation does not need it. Here Today is (supposed to be) an input whereas Names is (supposed to be) an output. This distinction is enforced by the user when the specification is executed. remind does not transition to a new state; it just outputs information. This is specified by not including the next-state variables in the head of the clause.
One more operation of the birthday book specification is given in A.
In summary, the specification is given by a declarative, abstract, logic description—similar to those written in B and Z. Sets, binary relations, functions and their operators are the main building blocks of the language. Properties (invariants) and operations are given in the same and only language.
3.1 Other elements of the language
The state machine specification language also supports parameters, axioms and user-defined theorems (Rossi and Cristiá,, 2025, Section 13.1.1). Parameters play the role of machine parameters and constants in B specifications and the role of variables declared in axiomatic definitions in Z. That is, parameters serve to declare the existence of some (global) values accessible to invariants and operations, but they cannot be changed by operations—i.e., there is no next-state value for a parameter. Axioms are used to state properties of parameters. User-defined theorems are used to state properties that can be deduced from axioms, invariants, operations or theorems that have already been declared.
3.2 Typing state machines
So far we have paid no attention to types in the specification. However, users can, but are not forced to, add typing information to each predicate (invariants, operations, etc.) declared in the specification to avoid some classes of errors. Here we type remind just as an example.
operation(remind).dec_p_type(rel(name,date),date,set(name))).    % type declarationremind(Birthday,Today,Names) :-  rres(Birthday,{Today},M) & dom(M,Names) & dec(M,rel(name,date)).The dec_p_type declaration provides the type for remind. The term remind(rel(name,date),date,set(name)) declares the type of each argument of the operation. For instance, rel(name,date) is the type of Birthday, date is the type of Today, and so on. The type of M is given by an explicit type declaration, dec(M,rel(name,date)). In turn, name and date are basic or uninterpreted types; rel(name,date) is the type of all binary relations with domain in name and range in date; and set(name) is the type of all sets whose elements are of type name. If t is a basic type then its elements are of the form t:$\langle elem\rangle$, where $elem$ is any Prolog atom or natural number.
4 Machine execution
As we have shown in Section 2, $\{log\}$ predicates are both, programs and specifications. State machines defined as in the previous section are no exception. Then, the operations of the specification of the birthday book can be executed as if they were the routines of some program. As $\{log\}$ programs are normally much less efficient than Prolog programs, we see them as functional prototypes. Running functional scenarios on a prototype helps, for example, to uncover possible mistakes in the specification early on, to analyze complex features, to analyze interactions among the operations, novice users to check if they have written the right predicates, etc. Functional scenarios can be executed directly from the $\{log\}$ interpreter in two different ways: by using the Next environment (Section 4.1), and by using symbolic execution (Section 4.2).
4.1 The Next environment
Next is a component tightly integrated with $\{log\}$ that simplifies the execution and analysis of deterministic, fully instantiated functional scenarios. Although Next is less general than the approach shown in the next section, it considerably ease the job of users. For example, we can add a person’s name and birth date when the system is in its initial state, as follows:
{log}=> initial >> addBirthday(Name:alice,Date:may24).  Known = {alice}, Birthday = {[alice,may24]}initial refers to the predicate declared as such in the specification; >> (read then) imposes a sequential order in the execution. In this way, $\{log\}$ automatically fetches state variables when calling an operation and chains the after-state of an operation with the before-state of the next one. Users need to indicate values only for the input variables (e.g., Name:alice). With Next is easy to add two persons to the system:
{log}=> initial >>        addBirthday(Name:[[alice,bob]],Date:[[may24,nov05]]).  Known = {alice,bob},  Birthday = {[alice,may24],[bob,nov05]}
The execution trace of the scenario can be analyzed by enclosing the initial state between square brackets:
{log}=> [initial] >>        addBirthday(Name:[[alice,bob]],Date:[[may24,nov05]]).  Known = {},  Birthday = {}    ----> addBirthday(Name:alice,Date:may24)  Known = {alice},  Birthday = {[alice,may24]}    ----> addBirthday(Name:bob,Date:nov05)  Known = {alice,bob},  Birthday = {[alice,may24],[bob,nov05]}
More operations can be called by using >>:
{log}=> [initial] >>        addBirthday(Name:[[alice,bob]],Date:[[may24,nov05]]) >>        remind(Today:may24,Names).
Next also performs invariant checking after each step of an execution. Users can indicate the invariants to be checked by appending a list with their names to the initial state. For example: [initial]:[domBirthday] would make Next to check domBirthday after each step of a scenario. If an invariant is not satisfied after a given step, then Next informs the user of the situation. This, in turn, provides valuable information on the correctness of the specification.
4.2 Symbolic executions
$\{log\}$ can perform symbolic executions too (King,, 1976). That is, we can execute the prototype by providing some variables instead of constants as inputs—or a mixture of both. In this way we can draw more general conclusions about the specification. In this case, though, we cannot use Next because it requires variables involved in the scenario to be completely instantiated. Actually, in symbolic executions we need to analyze the constraints returned by $\{log\}$. The following scenario analyses how remind responds to a call where Birthday has at least two equal dates in its range:
{log}=> pfun({[X,Y],[Z,Y]/B}) & remind({[X,Y],[Z,Y]/B},Y,Names).  Names = {X,Z/_N1}  Constraint: X neq Z, dom(_N5,_N1), un(_N4,_N5,B), ran(_N5,_N3),              subset(_N3,{Y}), ...We have deliberately cut out the list of constraints returned by $\{log\}$ to simplify the exposition. pfun({[X,Y],[Z,Y]/B}) restricts the set to be a function, thus forcing X neq Z. Then, the names returned by remind includes X and Z, plus some set _N1. The constraints state that _N1 is the domain of _N5, a subset of B, whose range contains only Y. That is, Names contains the right names and nothing else. Another scenario that can be analyzed is given by Birthday being a constant function:
pfun({[X,Y],[Z,Y]/B}) & ran(B,R) & subset(R,{Y}) &remind({[X,Y],[Z,Y]/B},W,Names).Note that in this case W, not Y, is the third argument passed in to remind. We analyze only two solutions. The first one is the following:
W = Y,  Names = {X,Z/N1}Constraint: ran(B,R), subset(R,{Y}), dom(N5,N1), un(N4,N5,B),            ran(N5,N3), subset(N3,{Y}), ran(N4,N2), Y nin N2, ...As can be seen, in this solution W = Y and Names includes X and Z plus some set N1. As above, N1 is the domain of N5, a subset of B, whose range contains only Y. Note that the rest of B, N4, has a range not containing Y. That is, N4 and N5 are a partition of B. The second solution is the following:
B = {[X,Y],[Z,Y]/N2},  R = {Y/N1}Constraint: Y neq W, dom(N5,Names), un(N5,N6,N2), ran(N2,N1),            subset(N1,{Y}), ran(N5,N4), subset(N4,{W}), ...In this case, Y neq W which, along with the fact that Y is the only element in the range of Birthday, implies that Names is the empty set. Names is the domain of N5 which in turn is a subset of N2 with a range containing only W. Then, Names is necessarily empty. This is confirmed by running the following:
{log}=> groundsol.{log}=> pfun({[X,Y],[Z,Y]/B}) & ran(B,R) & subset(R,{Y}) &        remind({[X,Y],[Z,Y]/B},W,Names) & W neq Y.  X = n0,  Y = n2,  Z = n1,  B = {},  R = {},  W = n3,  Names = {}
Analysis such as these increase our confidence on the correctness of the specification perhaps more than executions where inputs are constants because variables cover all possible constant values.
5 Machine consistency
One of the advantages of having a state machine language in $\{log\}$ is that a collection of standard VC can be automatically generated. Once discharged, these VC ensures the specification is consistent. Then, we have implemented a Verification Condition Generator (VCG) as a $\{log\}$ component. In this section we show how the VCG works.
Once users are done with executing the prototype, they can call the VCG which checks several well-formedness conditions about the structure of the specification and in that case generates a file containing $\{log\}$ code encoding the VC as well as code to discharge them. Some of the VC generated by the VCG are the following—the examples correspond to the birthday book specification.
  1. 1.
    The initial state satisfies each and every invariant. For instance:
    bbInit_sat_domBirthday :-  bbInit(Known,Birthday) & domBirthday(Known,Birthday).
  2. 2.
    Each operation is satisfiable and can change the state. For example:
    addBirthday_is_sat :-  addBirthday(Known,Birthday,Name,Date,Known_,Birthday_) &  [Known,Birthday] neq [Known_,Birthday_].If the operation does not change state variables, then the VC checks satisfiability of the operation.
  3. 3.
    Invariance lemmas: each operation preservers each and every invariant. For example888addBirthday_pi_birthdayFun stands for addBirthday preserves invariant birthdayFun.:
    addBirthday_pi_birthdayFun :-  neg(    birthdayFun(Birthday) &          addBirthday(Known,Birthday,Name,Date,Known_,Birthday_)          implies    birthdayFun(Birthday_)    ).
More VC are generated when the specification uses axioms and user-defined theorems—see $\{log\}$ user’s manual for more details (Rossi and Cristiá,, 2025, Section 13.6). The most important VC are the invariance lemmas. However, if operations or invariants are unsatisfiable, then invariance lemmas will trivially hold, thus, the VCG also generates the first two VC. Note that the first two classes of VC are satisfiability proofs whereas the third one entails unsatisfiability proofs. In general $\{log\}$ will be very fast at discharging the first two classes of VC and slower when discharging the third one.
5.1 Discharging verification conditions
After loading the file generated by the VCG, users can execute a small Prolog program (called check_vcs) that will attempt to discharge one VC after the other by passing in them to $\{log\}$. In other words, $\{log\}$ is used as an automated theorem prover. Recall that each VC is a $\{log\}$ formula written in terms of the predicates defined by the user whose bodies are, again, $\{log\}$ formulas written in terms of constraints. Users are informed about the status of each VC:
  • Ok: the VC has been successfully discharged.
  • Timeout: $\{log\}$ attempted to discharge the VC but it run out of time.
  • Error: $\{log\}$ found a counterexample contradicting the VC.
Here we analyze the second situation, whereas the third one is analyzed in Section 5.2. $\{log\}$ can timeout when solving a VC simply because many decision procedures for set theory are NP-complete, although in practice they work in many cases [^9] In $\{log\}$ most decision procedures are based on set unification which has been proved to be NP-complete (Dovier et al.,, 2006).. When a VC times out, we do not know whether the VC is indeed a theorem of the specification or not. Then, we should help $\{log\}$ to assist us to find out whether the VC is provable or not. To this end, users can rerun the VC by calling check_vcs with a couple of arguments. These arguments are passed in to $\{log\}$ making it to slightly change the solving algorithm as follows.
  • Timeout. We can simply extend the default timeout to see if $\{log\}$ needs some more time to discharge the VC.
  • Execution options. This is the more promising course of action as execution options influence the constraint solving algorithm. Influence means that the algorithm can take shorter to solve a formula but it also can take longer. It is hard to predict what execution options will have a positive impact in solving a particular formula. However there are some guidelines that in general produce good results. Here we will show a few execution options—see $\{log\}$ user’s manual for more details (Rossi and Cristiá,, 2025, Section 11.1).
    • subset_unify. Implements set equality as a double set inclusion instead of implementing it by exploiting set unification (Dovier et al.,, 2006). Note that set equality is pervasive in $\{log\}$ formulas even if the formula does not use it explicitly.
    • un_fe. Implements $un(A,B,C)$ in terms of RUQ (Section 2.4). In this way $\{log\}$ produces two solutions instead of six when solving $un(A,B,\{X/C\})$, with $A$ and $B$ variables. These two solutions encode the standard proof of $x\in A\cup B$. Union is not as pervasive as equality but is one the most used constraints during the low-level stages of the algorithm.
    • noirules. By default $\{log\}$ applies some (optional) inference rules while the goal is processed. noirules deactivates the application of these inference rules.
    • strategy(ordered). Changes the order in which atoms are processed. It has shown to impact proofs of the form $p\land(q\lor r)\land s$. Note that invariance lemmas have this form, indeed: $\lnot(Inv\land Op\implies Inv^{\prime})\equiv Inv\land Op\land\lnot Inv^{\prime}$, where $Op$ is usually a disjunction—e.g., addBirthday.
    • tryp(prover_all). This is the most powerful option. It attempts to solve the VC by running it in multiple (operating system) threads. Each thread runs the VC under some combination of execution options. As soon as one thread terminates, the whole computation terminates as well. In this way, the net execution time will tend to be the time needed by the thread running the best combination of execution options for that VC.
    Therefore, users can analyze the VC to determine what execution options would be useful for that particular formula. If they have access to a large computer tryp(prover_all) is the best course of action as it tries all possible combinations at once. According to our experiments there is always a combination of execution options that discharges the VC in a reasonable time. At the same time, exploring execution options is far less cumbersome than attempting an interactive proof.
5.2 Analyzing undischarged verification conditions
$\{log\}$ may not be able to discharge a VC for a few reasons: (i) If a VC falls outside the decision procedures implemented in $\{log\}$, then the tool will be unable to discharge it; there is nothing to do in this case—see Section 5.3. (ii) If the specification is wrong (e.g., a precondition is missing, an invariant is too strong, etc.) then it may be impossible to prove some VC. (iii) Finally, a verification condition may require more hypotheses. Let’s analyze these last two situations.
Every time the status of a VC is Error, $\{log\}$ saves a counterexample contradicting the VC. These counterexamples can be very helpful in finding out why the proof failed. There are two kinds of counterexamples: abstract, which may include free variables; and ground, which do not contain free variables. For example, when $\{log\}$ attempts to discharge the VC of the birthday book specification, the VC named addBirthday_pi_birthdayFun fails. We can see the ground counterexample as follows101010vcgce stands for verification condition ground counter example.:
{log}=> vcgce(addBirthday_pi_birthdayFun).  Birthday = {[n2,n1]}, Known = {}, Name = n2, Date = n0,  Known_ = {n2}, Birthday_ = {[n2,n0],[n2,n1]}The reason for the failure is in this counterexample. As can be seen, Birthday = {[n2,n1]} whereas Known = {} thus contradicting invariant domBirthday which states dom(Birthday,Known). That is, the counterexample says that Birthday is not empty while Known, its domain according to domBirthday, is empty. Therefore, the proof failed because it lacks domBirthday as an hypothesis. Once the user adds domBirthday as an hypothesis to addBirthday_pi_birthdayFun $\{log\}$ proves the VC immediately. Although in this case the cause of the problem is a missing precondition, the same analysis is conducted when the cause ends up being an error in the specification.
[…]
Aligned with the way $\{log\}$ generates VC, is the decision of splitting the invariant of the birthday book specification into two predicates (domBirthday and birthdayFun). Indeed, these predicates could be defined as a single invariant. However, this strategy may hinder the automated proof process as it is like having all the invariants as hypotheses.
We have applied this methodology, for example, to analyze Chinese wall security properties (Capozucca et al.,, 2024), where we have to add 12 hypotheses in 5 VC out of 35, that is, 30 VC need no human intervention; and to verify a model of a landing gear system (Cristiá and Rossi, 2024a, ), where 60% of the VC where discharged in the first run whereas the vast majority of the remaining proofs needed only one hypothesis.
In summary, $\{log\}$ is used to automatically discharge VC and to help users to find why a VC could not be discharged.
5.3 What verification conditions can be discharged?
In the previous section we mentioned that $\{log\}$ will be unable to discharge a VC if it falls outside the decision procedures implemented in it. A formal and precise description of the boundaries of the decidable fragments available in $\{log\}$ has been given in our previous work—see at the beginning of Section 2. In this section we give an informal account of the decision procedures implemented in $\{log\}$. Figure 1 helps us in this task. If theory A is above theory B it means that A inherits the signature of B.
  • LIA stands for linear integer arithmetic. $\{log\}$ uses an external implementation of a decision procedure for LIA.
  • SET is the Boolean algebra of sets. That is, a theory including hereditarily finite sets, equality, union, intersection, difference, set membership and their negations (Dovier et al.,, 2000).
  • CARD is SET extended with cardinality and combined with LIA (Cristiá and Rossi,, 2023). That is, variables denoting set cardinality can only participate in linear integer constraints. CARD cannot be combined with RIS or RA.
  • FII is CARD extended with finite integer intervals (Cristiá and Rossi, 2024c, ). As with CARD interval limits can only participate in linear integer constraints.
  • RIS is SET extended with RIS—Section 2.2. Considering RIS of the form ris(X in A,$\phi$), the theory is decidable as long as $\phi$ belongs to a decidable theory supported by $\{log\}$. When more complex RIS are used, the condition for decidability is more involved (Cristiá and Rossi, 2021b, ).
  • RA (relation algebra) is SET extended with composition, converse and the identity relation. The condition for decidability is rather involved (Cristiá and Rossi,, 2020); we give a simplified version. RA is decidable for the class of formulas not containing constraints of the form comp(R,S,R) or comp(S,R,R). As many relational constrains are defined in terms of comp (e.g., relational image), a formula not containing comp may still belong to the undecidable fragment—for example, some formulas containing (simultaneously) domain and range fall outside of the decidable fragment (Cantone and Longo,, 2014).
LIASETCARDRISRAFIIdecidableundecidableLegend
Figure 1: The stack of theories dealt with by $\{log\}$
The theory of restricted quantifiers, noted RQ, is not depicted in Figure 1. RQ is parametric w.r.t. some quantifier-free, decidable theory (Cristiá and Rossi, 2024d, )—for instance RQ could be configured with CARD or LIA or even with the decidable fragment of RA. However, the parameter theory must verify the following condition: if its solver, as a result of solving a formula, returns a conjunction of atoms including set variables, this conjunction must be satisfiable by substituting set variables by the empty set. This is not the case of CARD and FII. The decidability of RQ is as follows. Let $\{\dots/A\}$ be the domain of a RQ where $A$ is a variable. We say that $A$ is the domain variable of the RQ [^12] Clearly, if the domain of a RQ is just a variable then this is the domain variable.. The innermost formula of a RQ is called its quantifier-free formula. For instance, X is V + 1 is the quantifier-free formula of the nested RQ given in (2.4). The decidable fragment of RQ is characterized by formulas including foreach and exists with a quantifier-free formula belonging to a decidable fragment and meeting at least one of the following conditions.
  1. 1.
    The formula contains only exists.
  2. 2.
    The formula contains only foreach, and none of the domain variables are used in the quantifier-free formula.
  3. 3.
    The formula contains nested RQ of the following form where foreach verify (2):
    exists([$X_{1}$ in $A_{1}$,...,$X_{n}$ in $A_{n}$],  foreach([$Y_{1}$ in $B_{1}$,...,$Y_{m}$ in $B_{m}$], $\phi$))
  4. 4.
    The formula contains foreach and exists but does not belong to the above class (i.e., some exists occur after some foreach). In this case the condition for decidability is as follows: (i) no domain variable of a foreach is used in the quantifier-free formula; and (ii) no exists occurring after a foreach share the same domain variable. For instance, the following formula do not verify 4(ii): foreach(X in {W / A}, exists(Y in {V / A}, $\phi$)).
If $\{log\}$ is called on a formula not meeting the conditions for the theories depicted in Figure 1 nor the conditions for RQ, it will most likely run forever although it may terminate in some cases. Then, in general, calling $\{log\}$ on such a formula will be rather harmless. There is, however, an exception. If the formula combines CARD with RIS, RA or RQ, then the answer can be unsound.
6 Test case generation
As we have said, the state machines specified in $\{log\}$ are considered functional prototypes. These specifications will likely be implemented in some other programming language. These implementations should be tested. Model-based testing (MBT) (Utting and Legeard,, 2006) indicates that implementations can be tested with test cases generated from their specifications. $\{log\}$ includes an implementation of the Test Template Framework (TTF) (Stocks and Carrington,, 1996), a MBT method originally developed for the Z notation. The TTF was first implemented in Java for Z specifications (Cristiá et al.,, 2014) and later connected to $\{log\}$ (Cristiá et al.,, 2013). In that implementation users are required to write a Z specification part of which is translated into $\{log\}$ and the results computed by $\{log\}$ are translated back to Z. Now the TTF has been implemented as a $\{log\}$ component that generates test cases from state machines specified in $\{log\}$. As a result we have a much simpler and cleaner tool seamlessly integrated with $\{log\}$.
We will use the birthday book specification to show how the TTF works in $\{log\}$. Before applying the TTF on a specification, the specification has to be typechecked and the VCG has to be called on that specification. The TTF uses information about the state machine collected by the typechecker and the VCG. Test cases are generated for each operation in a state machine. In this example we will generate test cases for addBirthday. Test case generation proceeds by applying so-called testing tactics to the selected operation. What tactics are applied depends on the logical structure and mathematical elements used in the operation, as well as on the testing goals determined by the testing team. In $\{log\}$ the first tactic to be applied is called disjunctive normal form (DNF). The command is as follows:
[…]
addBirthday_vis    addBirthday_dnf_1    addBirthday_dnf_2addBirthday_vis is the root of the tree while addBirthday_dnf_1 and addBirthday_dnf_2 are its children—indentation is used to depict tree levels. Nodes in the testing tree are called test specifications. Each test specification is a conjunction of $\{log\}$ constraints depending on the before-state and input variables of the operation. Although we have not passed in state variables to applydnf, it uses the structure of the specification to get them. This tactic writes the precondition of the operation into DNF and generates a test specification for each term of the DNF. The following are the test specifications generated so far:
addBirthday_dnf_1(Name,Date,Known,Birthday) :- Name nin Known.addBirthday_dnf_2(Name,Date,Known,Birthday) :- Name in Known.
It would be possible to generate test cases from this testing tree but we will apply one more tactic in order to show how $\{log\}$ is used by the TTF more thoroughly. After DNF, tactics are applied to one or more test specifications of the testing tree. Then, we apply a second tactic known as standard partitions (SP) to the addBirthday_dnf_1 test specification. The command is as follows:
{log}=> applysp(addBirthday_dnf_1,un(Known,{Name},Known_)).The first argument is the test specification and the second one is a constraint present in the operation—see addBirthdayOk in Section 3. SP defines partitions for key mathematical operators such as union, intersection, overriding, etc. Figure 2 shows the standard partition for union, intersection and set difference. The partition indicates conjunctions of conditions depending on the arguments of the operator. The rationale behind SP is that set and relational operators have non trivial implementations that deserve a thorough testing. The SP tactic generates a test specification for each conjunction defined in the partition. Then, the above command generates the following testing tree.
addBirthday_vis    addBirthday_dnf_1        addBirthday_sp_11        addBirthday_sp_12        .................   % some test specifications not shown        addBirthday_sp_17        addBirthday_sp_18    addBirthday_dnf_2addBirthday_dnf_1 is partitioned into eight test specifications due to the eight conjunctions in the standard partition for set union. Two sample test specifications are the following:
addBirthday_sp_11(Name,Date,Known,Birthday) :-  Name nin Known & Known = {} & {Name} = {}.addBirthday_sp_14(Name,Date,Known,Birthday) :-  Name nin Known & Known neq {} & {Name} neq {} & disj(Known,{Name}).Note that each test specification includes the predicate of addBirthday_dnf_1. Besides, note that the arguments of the partition ($S$ and $T$) are substituted by the arguments of the constraint passed in to the command ({Name} and Known).
$S=\emptyset,T=\emptyset$ $S\neq\emptyset,T\neq\emptyset,S\subset T$
$S=\emptyset,T\neq\emptyset$ $S\neq\emptyset,T\neq\emptyset,T\subset S$
$S\neq\emptyset,T=\emptyset$ $S\neq\emptyset,T\neq\emptyset,T=S$
$S\neq\emptyset,T\neq\emptyset,S\cap T=\emptyset$ $S\neq\emptyset,T\neq\emptyset,S\cap T\neq\emptyset,S\not\subseteq T,T\not\subseteq S,S\neq T$
Figure 2: Standard partition for $S\cup T$, $S\cap T$ and $S\setminus T$
Finally, observe that addBirthday_sp_11 is unsatisfiable. This is a usual situation in the TTF. Given that it is impossible to generate test cases from unsatisfiable test specifications, they must be pruned from the testing tree. Detecting unsatisfiable set formulas is one of the main $\{log\}$’s capabilities. Hence, the TTF provides a command, namely prunett, that iterates over all the leaves of a testing tree and asks $\{log\}$ whether or not they are unsatisfiable. If $\{log\}$ finds an unsatisfiable test specification, prunett eliminates it from the testing tree. The resulting testing tree after calling prunett is the following:
addBirthday_vis    addBirthday_dnf_1        addBirthday_sp_12        addBirthday_sp_14    addBirthday_dnf_2
We could apply more testing tactics but that would not add anything substantial to the understanding of the $\{log\}$ framework. The final step in the TTF is generating one test case for each leaf in the testing tree. In this context a test case is an assignment of values to the before-state and input variables appearing in the test specification. In other words, a test case is a solution to some test specification. Again, $\{log\}$ is good at finding solutions to set formulas. Then, the TTF provides a command, namely gentc, that iterates over all the leaves of a testing tree and asks $\{log\}$ for a ground solution to each of them. The test cases generated by gentc are the following:
addBirthday_tc_12(Name,Date,Known,Birthday) :-  Name = name:n0 & Known = {}.addBirthday_tc_14(Name,Date,Known,Birthday) :-  Name = name:n0 & Known = {name:n1}.addBirthday_tc_2(Name,Date,Known,Birthday) :-  Name = name:n0 & Known = {name:n0}.Observe that Birthday has not been bounded to some value. This is so because no test specification includes a condition on that variable. Concerning these test cases, Birthday can assume any value. Applying SP on the postcondition un(Birthday,{[Name,Date]},Birthday_) as a third tactic would generate test cases with different values for Birthday combined with the values of these test cases. B presents the application of this tactic and the test cases so generated.
In summary, the TTF component uses $\{log\}$ as a back-end solver for automating two crucial but annoying, error-prone tasks: pruning test specifications and test case generation. Users are left with the more creative task of deciding what tactics will generate the best test suite. The current implementation of the TTF provides a total of six testing tactics; SP supports nine set and relational operators but can be easily extended by users—see $\{log\}$ user’s manual for more details (Rossi and Cristiá,, 2025, Section 13.7).
7 Concluding remarks
After adding a few components we were able to turn the CLP language and solver $\{log\}$ into a formal verification tool. It seamlessly integrates implementations of some sought-after theoretical results such as symbolic execution, automated proofs and the formula-program duality. From the practical side, $\{log\}$ now offers a state machine specification language based on FOL and set theory, an environment that simplifies the analysis of functional scenarios, a verification condition generator that fed proof obligations into $\{log\}$, and a model-based test case generator that uses $\{log\}$ as the back-end solver. There is just a single language for programs, specifications and properties; the same solver runs functional scenarios, proves properties of the specification and generates test cases.
Nonetheless, $\{log\}$ needs more functions to get closer to tools such as Agda and Dafny. To name some of them: 1. Now users have to manually add hypotheses to VC which is cumbersome and error-prone. A user command could perform safety checks before the hypothesis is added. 2. Currently, $\{log\}$ does not support refinement proofs, at least not in a structured fashion.
More advanced features should also be analyzed. For instance, first-order automated reasoning, such as the one performed by the E-prover (Schulz et al.,, 2019), may help in finding missing hypotheses. That is, $\{log\}$ could rest in E-prover to select a smaller set of promising hypotheses and then try to discharge the VC using its own engine. An even more challenging issue is to address the proofs of VC falling outside of the decision procedures available in $\{log\}$. An obvious solution would be to rest on an external interactive theorem prover (ITP). But using an external ITP brings in the problem of translating proof terms back and forth, thus introducing new sources of errors. On the other hand, even such proofs will have subgoals fitting in the decidable fragments where $\{log\}$ can be handy. Therefore, we have developed a simple prototype ITP on top of $\{log\}$ where users can call it whenever they have a subgoal fitting in some decidable fragment (Cristiá et al.,, 2022). Nevertheless, this idea needs considerably more work.
Competing interests: The authors declare none
Appendix A One more operation of the birthday book
The remaining operation of the birthday book shows the birthday of a given person.
findBirthdayOk(Known,Birthday,Name,Date) :-  Name in Known & applyTo(Birthday,Name,Date).
notAFriend(Known,Name) :- Name nin Known.
operation(findBirthday).findBirthday(Known,Birthday,Name,Date) :-  findBirthdayOk(Known,Birthday,Name,Date) or notAFriend(Known,Name).
Appendix B Extended application of the TTF to the addBirthday operation
The following TTF script has been applied to the addBirthday operation of the birthday book specification.
applydnf(addBirthday(Name,Date)).
applysp(addBirthday_dnf_1,un(Known,{Name},Known_)).
prunett.
applysp(addBirthday_dnf_1,un(Birthday,{[Name,Date]},Birthday_)).
prunett.
gentc.
In this way the following nine test cases are generated.
addBirthday_tc_122(Name,Date,Known,Birthday) :-
  Name=name:n1 & Known={} & Birthday={} & Date=date:n0.
addBirthday_tc_124(Name,Date,Known,Birthday) :-
  Name=name:n1 & Known={} & Birthday={[name:n0,date:n0]} & Date=date:n2.
addBirthday_tc_126(Name,Date,Known,Birthday) :-
  Name=name:n1 & Known={} &
  Birthday={[name:n1,date:n2],[name:n0,date:n0]} &
  Date=date:n2.
addBirthday_tc_127(Name,Date,Known,Birthday) :-
  Name=name:n1 & Known={} & Birthday={[name:n1,date:n0]} & Date=date:n0.
addBirthday_tc_142(Name,Date,Known,Birthday) :-
  Name=name:n0 & Known={name:n1} & Birthday={} & Date=date:n2.
addBirthday_tc_144(Name,Date,Known,Birthday) :-
  Name=name:n0 & Known={name:n1} &
  Birthday={[name:n2,date:n2]} &
  Date=date:n3.
addBirthday_tc_146(Name,Date,Known,Birthday) :-
  Name=name:n0 & Known={name:n1} &
  Birthday={[name:n0,date:n3],[name:n2,date:n2]} &
  Date=date:n3.
addBirthday_tc_147(Name,Date,Known,Birthday) :-
  Name=name:n0 & Known={name:n1} &
  Birthday={[name:n0,date:n2]} &
  Date=date:n2.
addBirthday_tc_2(Name,Date,Known,Birthday) :-
  Name=name:n0 & Known={name:n0}.